OpenStudy (fatimab2000):

will give medal !!!!!! 1. The average annual salary of the employees of a company in the year 2005 was $70,000. It increased by the same factor each year and in 2006, the average annual salary was $82,000. Let f(x) represent the average annual salary, in thousand dollars, after x years since 2005. Which of the following best represents the relationship between x and f(x)? f(x) = 70(1.17)x f(x) = 82(1.17)x f(x) = 70(2.2)x f(x) = 82(2.2)x

OpenStudy (fatimab2000):

2.The price of gold has increased by 35% per year from 2000. In the year 2000, Harry bought a gold ring for $590. Which of the following functions f(x) can be used to represent the price of the ring x years after 2000? f(x) = 590(1.35)x f(x) = 590(0.65)x f(x) = 35(0.41)x f(x) = 35(1.59)x

OpenStudy (fatimab2000):

@Michele_Laino

OpenStudy (fatimab2000):

@ganeshie8

OpenStudy (fatimab2000):

@imqwerty

OpenStudy (just_one_last_goodbye):

Need help? ^_^

OpenStudy (just_one_last_goodbye):

"You need to find the factor of increase from 2005 to 2006. Since you are given the amounts 70,000 and 82,000 it should be fairly simple to find. this will be you rate or r value."

OpenStudy (just_one_last_goodbye):

820000/70000 = 1.17

OpenStudy (just_one_last_goodbye):

P = 70

OpenStudy (gabebae):

1. F(x) = 82(1.17)^x 2. F(x) = 590(1.35)^x principal *interest rate^time

OpenStudy (just_one_last_goodbye):

@GabeBae no direct answers

OpenStudy (gabebae):

Sorry?

OpenStudy (just_one_last_goodbye):

http://openstudy.com/code-of-conduct